Punch List 

Currently Broken 

As-built drawings 
Need them after absolutely all changes/fixes finished. 
Is it possible to also get machine-readable stuff in addition to paper? 

if send reset, it hangs. speculation: iff Y/C card is in a specific slot? 
Yes. Only hang if reset the YC level while the YC card is in place. Restting 
other levels, or pulling the YC card prevent the system hang. 

DPS Remote 
Control of both Beta internal TBCs from DPS remote station 

Possible Problems / Needs Testing 

Verify the A60 SDI path since we got new cards 

Verify the A20 

Verify signal through DPS for gain and other quality 

Hires RGB Signal is rather munged when going through Covid DAs.
I couldnt verify this

conference room NTSC and audio. Snader tested, we should doublecheck.


I/O panel. Snader tested, we should doublecheck.

CAV_RTR - is this the Faroudja??
video comes up green or distorted depending on encoder switch settings; may be problem with CAV->NTSC encoder

robustness of the ADC6801, since the DEC6801 was less than stunning.
Needs to be dealt with (not Snader's problem) 
  •  Hi-8 can only have ntsc or svideo, not both plugged in. A possible solution is a svideo switchbox. that we mount on top of the Hi-8. A manual one should be cheap and then we can switch svideo and ntsc (with svideo is non connect) on the front of the hi-8. We could also ignore the issue.
  • Calibrate For Gain
    • Lyon Lamb
    • FORA
    • Faroudja
    • Beta
    • FS vs. bars?
  • Audio Black - should this be disconnect or something real?s
  • Get Faroudja fixed
    Runs way too hot, input and outputs need alignment. Send to Faroudja in Sunnyvale to do this. Not Snader's problem.


    We have to do full documentation. Sigh.
    Need to set up keyboard/mouse in conference room.
    Serial control
    Infomural inputs
    Need cables under floor. Should be properly labelled!
    PC input to patch panel
    Analog on the Galileo
    Analog input on the galileo seems to be missing a chroma channel. Moving CAV to input 2, (which neccesitates moving NTSC to input 1) fixes the problem. Our galileo is just broken on that input I think.
    Serial on Galileo
    I got it working. Its an SGI problem, not snader. This sequence of steps can break/fix SDI output on glim. FIX SDI: Connect SDI input cable to input #2l. Bring up a digital video input window. Close Window. Use SDI output. (I've done this twice) Break SDI: A bunch of stuff I can't duplicate (It started broken, and I broke it once.) We need to leave the input on #1 though since input #2 and the output can't be active at the same time. There are several options. Just move the cable to fix. Mirror the #1 and #2 inputs. Have a different input from the rack to #2. (incidentally the galileo is our only real time two video stream device so why don't we have that extra input connected anyway.)
    SITN through IDEN by default??
    If we did this by default, then SITN signal would be genlocked. Hmm. But would be affected by wierd gain
    SadButTrue: IDEN TBC dub does not increase quality
    Just treat the IDEN as a TBC. Ignore the fact that there's also a dub connection from the play UMatic. Just route through the matrix.

    Tested/Fixed (used to be above) 

    a CCD camera tested and works. jed changed it from RGB to YUV/CAV 

    lightning inputs and outputs - tested but not connected - You need a sync adapter to get VGAs two syncs into something useable. I used an adapter that puts sync on green. 

    LyonLamb (yuv) -> Beta (sp and digi) -> monitor (yuv)
    Shifted by 1/4 vertically but stable.

    Is this fixed??? Seems to be 
    SITN scrolling (no sync) through RGB Decoder
    SITN isn't genlocked to the rest of the system so of course this happens. Other NTSC inputs do fine. Using a FrameSync before the RGB Decoder fixes the problem.
    This scrolling has to do with the sync on the monitor, not the decoder. Need to choose
    either composite or compnent sync on the front panel.
    Audio levels right??
    Ive been told that audio levels have been adjusted. However looking at the mixer, the left and right tone inputs are not equal. I went looking for the right place to adjust this, but couldn't find it. I'm happy to make the adjustment if someone tells me which gain knob to mess with.
    Adjust Faroudja
    Broken: Sitn -> Frame Sync (ntsc) -> Lietch ntsc-to-cav -> Faroudja -> Monitor
    Shifted to the right (like I showed paul from the digi-beta before) It can't be the Faroudja since item 5 works. And it can't be the Leitch since item 6 works, so Im stumped. They dont work together. Why?
    This was explained by the fact that the DEC6801 destroys frame sync, and the faroudja
    requires it.
    Telect -
    audio frame was miswired, now inputs 7/8 on mixer and other stuff works?
    Fixed : LyonLamb (yuv) ->Faroudja -> Monitor (ntsc)
    The LyonLamb and Faroudja are both supposed to be genlocked, but this scrolls. Why? 

    Answer: Black burst genlock was hooked up to wrong thing - was on sync in. jed moved it to composite in.
    Questions/Comments (videolab@graphics.stanford.edu)